Solve the following triangles:


13. A side and a diagonal of a parallelogram are 12 inches and 19 inches respectively. The angle between the diagonals, opposite the given side, is 124°. Find the length of the other diagonal and the length of the other side.

14. A lighthouse is 10 miles northeast of a dock. A ship leaves the dock at noon, and sails east at a speed of 12 miles an hour. At what time will it be 8 miles from the lighthouse?

15.A vertical pole 35 feet high, standing on sloping ground, is braced by a wire which extends from the top of the pole to a point on the ground 25 feet from the foot of the pole. If the pole. subtends an angle of 30° at the point where the wire reaches the ground, how long is the wire?

16.A tower 125 feet high stands on the side of a hill. At a point 240 feet from the foot of the tower, measured straight down the hill, the tower subtends an angle of 25°. What angle does the side of the hill make with the horizontal?


1
Expert's answer
2021-12-13T17:31:33-0500

(13)

Let b"=" Side adjacent to angle of "124\u00b0"

Using cosine rule


"12^2=9.5^2+b^2=2(b)(9.5) cos 124"

"b^2+10.62b-53.75=0"



"b=\frac{-10.62{^+_-}\sqrt{10.62^2+4(53.75)}}{2}"


"b=3.74" or "-14.36"

Length of the diagonal "=3.74\u00d72"

"=7.48inches"


(14)


Let b "=" distance sailed.

Using cosine Rule


"8^2=b^2+10^2-2×10×b\>cos\>45"

"b^2-14.14b+36=0"


"b=\frac{14.14{^+_-}\sqrt{14.14^2-4(36)}}{2}"

"=3.329\>or\>10.813"



Time taken "=\frac{3.329}{12}×60=16'39''"


Or "\frac{10.813}{12}×60=54'4''"


Time can be";"

"12:16'39''\>p.m"

"Or\>12:54'4''\>p.m"


(15)


Let b"=" length of the wire

Using cosine rule


"35^2=b^2+25^2-2×b×25\>cos\>30"


"b^2-43.30b-600=0"


"b=\frac{43.30{^+_-}\sqrt{43.3^2-4(-600)}}{2}"


"=54.34\>or\>-11.04"

Length of the wire "=54.34feet"



(16)


From a triangle 240 feet opposite to "\empty°" and 125 feet opposite to "25\u00b0"


"\therefore" "\frac{240}{Sin{\empty}}=\frac{125}{Sin25}"


"Sin{\empty}=\frac{240\>Sin25}{125}"


"{\empty}=54.24°"


Let "\theta""=" angle to horizontal

"\theta=90-(54.24+25)"

"=10.76\u00b0"